OpenStudy (anonymous):

Please help. Solve these equations write step by step if you can 1. x +3y -z=10 2x+5y+4z=57 3x -y+ 2z=15 2. x+ y +z= 53 x+2y+3z=105 x+3y+4z=134 3. 2x-3y+z=-2 4x-4y-3 =2 6x+y-4z =6 Thank you so much!

OpenStudy (cwrw238):

1. first multiply the first equation by -2 then add the resulting equation to the second equation this will eliminate the term in x and give you an equation in y and z next multiply the first equation by -3 then add to equation 3 this will give you another equation in y and z now solve these 2 equations in y and z by the elimination method or substitution method finally plug in the values of y and z into one of the original equation to find x

OpenStudy (cwrw238):

and you ok with that explanation? i'll do the first poart for you x + 3y - z = 10 multiply by -2:- -2x - 6y + 2z = -20 2x +5y + 4z = 57 add: -y + 6z = 37

OpenStudy (cwrw238):

ok?

OpenStudy (cwrw238):

r u there?

OpenStudy (anonymous):

Ok thanks, the steps were just confusing It is the same for #2?

OpenStudy (cwrw238):

basically yes if you multiply the first equation by -1 then add the second equation you will eliminate x